Find 3 consecutive odd integers such that twice the 2nd decreased by the 3rd is 3 less than twice the 1st
Found 2 solutions by user_dude2008, stanbon:
Answer by user_dude2008(1861)

2(x+2)-(x+4)=2x-3
2x+4-x-4=2x-3
x=2x-3
x-2x=-3
-x=-3
x=3 <-------------- First #
x+2=5 <-------------- Second #
x+4=7 <-------------- Third #

Answer by stanbon(57347)

Find 3 consecutive odd integers such that twice the 2nd decreased by the 3rd is 3 less than twice the 1st
-------------
1st: 2x-1
2nd: 2x+1
3rd: 2x+3
-------------------------
Equation 2(2x+1) - (2x+3) = 2(2x-1)-3
4x+2-2x-3 = 4x-2-3
2x-1 = 4x-5
2x=4
x = 2
----------
1st: 2x-1 = 3
2nd: 5
3rd: 7
------------------
Cheers,
Stan H.